# AnySR: Realizing Image Super-Resolution as Any-Scale, Any-Resource

### v1

@@ 4.4 Ablation Studies
Table 4: Ablations on alternations of AnySR including the “w/o ASE” variant removing any-scale enhancement (ASE) and the “w/o FI” variant replacing feature-interweaving (FI) with a simple concatenation. The denotes AnySR variants (through different subnets) and denotes AnySR-retrained version (through the largest network). Underlines indicates the best performance.
@@
Table 5: Ablations on the reset probability $p$. The denotes AnySR-retrained version (through the largest network). Underlines indicates the best performance.
@@
Any-Scale Enhancement. Any-scale enhancement constitutes one of the fundamental branches encompassed within our overarching research framework. By injecting, excavating, and ameliorating sufficient scale information, we realize customized handling for features at different scales. In order to validate the effectiveness of ASE, we train the network by removing this component, and the results are presented in Table 4. It is evident that the absence of ASE (i.e., “w/o ASE”) leads to a certain performance drop.
Feature-Interweaving. Feature-interweaving considers the non-uniformity of shared weights and overcomes the mutual weight influence across different scales by repeating and inserting scale pairs into features at regular intervals. We study the contribution of this mechanism by substituting the feature-interweaving fashion with a simple concatenation in earlier methods [10, 52]. By introducing feature-interweaving, a better performance is achieved in Table 4.
Reset Probability $p$. As mentioned in Sec. 3.2, we reset subnet $F_{t}$ to the entire network $F$ with a certain probability $p$ to retain the original ability. We conduct ablation experiments with different values of $p=\{0,0.2,0.4,0.6,0.8\}$. experimental groups. Note that $p=0$ indicates the removal of the reset mechanism. The results are presented in Table 5. A low $p$, in particular for $p=0$, leads to minimal impact on enhancing the performance of the entire network during inference. Hence, opting for $p=0.6$ or $0.8$ is deemed more appropriate. Nevertheless, it is observed that with $p=0.8$, the performance improvement for the entire network inference is slightly higher than $p=0.6$, and a higher probability incurs higher training costs. Considering the trade-off between the performance of the entire network inference and network training costs, we ultimately select $p=0.6$ as the experimental setting for the study presented in this paper.
5 Conclusion and Future Work

### v2

@@ IV-E Ablation Studies
TABLE VII: Ablation results of PSNR(dB) and LPIPS on alternations of AnySR including the “w/o ASE” variant removing any-scale enhancement (ASE) and the “w/o FI” variant replacing feature-interweaving (FI) with a simple concatenation. Underlines indicates the best performance.
@@
[…]
AnySR 0.041 0.092 0.138 0.072 0.164 0.227 0.112 0.230 0.304 0.048 0.120 0.176 0.016 0.046 0.078
AnySR (w/o ASE) 0.045 0.096 0.142 0.072 0.167 0.230 0.116 0.235 0.305 0.053 0.121 0.179 0.019 0.053 0.080
AnySR (w/o FI) 0.044 0.095 0.139 0.075 0.167 0.229 0.115 0.235 0.305 0.049 0.122 0.179 0.018 0.049 0.079
AnySR 0.040 0.093 0.138 0.071 0.163 0.227 0.112 0.230 0.304 0.047 0.119 0.176 0.016 0.046 0.078
AnySR (w/o ASE) 0.043 0.096 0.142 0.073 0.165 0.230 0.115 0.234 0.305 0.052 0.121 0.179 0.018 0.052 0.080
AnySR (w/o FI) 0.043 0.095 0.139 0.072 0.165 0.229 0.114 0.233 0.305 0.049 0.121 0.179 0.018 0.048 0.079
TABLE VIII: Ablation results of PSNR (dB) on the reset probability $p$. Underlines indicates the best performance.
@@
AnySR Variants
AnySR (p=0) 38.01 34.45 32.37 33.73 30.45 28.77 32.18 29.13 27.64 32.25 28.44 26.44 38.77 33.77 30.82
AnySR (p=0.2) 38.03 34.45 32.37 33.73 30.45 28.76 32.18 29.13 27.64 32.25 28.44 26.45 38.76 33.77 30.82
AnySR (p=0.4) 38.03 34.47 32.37 33.72 30.45 28.77 32.19 29.13 27.65 32.26 28.43 26.44 38.77 33.77 30.83
AnySR (p=0.6) 38.04 34.50 32.38 33.74 30.47 28.78 32.19 29.15 27.65 32.27 28.45 26.45 38.78 33.77 30.83
AnySR (p=0.8) 38.03 34.49 32.38 33.73 30.46 28.77 32.18 29.14 27.65 32.26 28.44 26.45 38.78 33.78 30.82
AnySR-Retrained Version
@@
Any-Scale Enhancement. Any-scale enhancement constitutes one of the fundamental branches encompassed within our overarching research framework. By injecting, excavating, and ameliorating sufficient scale information, we realize customized handling for features at different scales. To validate the effectiveness of ASE, we train the network by removing this component, and the results are presented in Table VII. It is evident that “w/o ASE” leads to a certain performance drop.
Feature-Interweaving. Feature-interweaving considers the non-uniformity of shared weights and overcomes the mutual weight influence across different scales by repeating and inserting scale pairs at regular intervals. We study this mechanism by substitute the feature-interweaving fashion with a simple concatenation in earlier methods [10, 53]. With feature-interweaving, a better performance is achieved in Table VII.
Reset Probability $p$. As mentioned in Sec. III-B, we reset subnet $F_{t}$ to the entire network $F$ with a certain probability $p$ to retain the original ability. We conduct ablation experiments with different values of $p=\{0,0.2,0.4,0.6,0.8\}$. Note that $p=0$ indicates the removal of the reset mechanism. The results are presented in Table VIII. A low $p$, in particular for $p=0$, leads to minimal impact on enhancing the performance of the entire network during inference. Hence, opting for $p=0.6$ or $0.8$ is deemed more appropriate. Nevertheless, it is observed that with $p=0.8$, the performance improvement for the entire network inference is slightly higher than $p=0.6$, and a higher probability incurs higher training costs. Considering the trade-off between the performance of the entire network inference and network training costs, we ultimately select $p=0.6$ as the experimental setting for the study presented in this paper.
V Conclusion and Future Work